OpenStudy (anonymous):

f(x) = 3 cos(3ln(x)) find f'(x)

OpenStudy (anonymous):

\[\eqalign{ &\frac{d}{dx}3\cos[3\ln(x)] \\ =&3\frac{d}{dx}\cos[3\ln(x)] \\ =&3\left[-\sin(3\ln x)\times3\times\frac{1}{x}\right] \\ =&-\frac{9\sin(3\ln x)}{x}}\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No prob man!